OpenStudy (anonymous):

How many liters of oxygen gas, at standard temperature and pressure, will react with 25.0 grams of magnesium metal? Show all of the work used to solve this problem. 2 Mg + O2 2 MgO I believe the equation to use is PV=nRT R= 0.0821 P= 1 atm T= 273Kelvin I'm pretty sure i have that so far correct? how do i figure out how many moles there are?

OpenStudy (anonymous):

and how do i figure out what V is??

OpenStudy (frostbite):

You are right to use the ideal gas equation (but take into account it is not going to be precise). The amount of moles, can be found using the relationship between mass, molecular mass and moles.

OpenStudy (frostbite):

How would you determine n(O2) in any other assignment knowing you have 25.0 grams of magnesium metal?

OpenStudy (anonymous):

uhm, i'm not sure? i would have to divide that by something, no? @Frostbite

OpenStudy (frostbite):

You are on the right track. Remember the equation? \[n=\frac{ m }{ M }\]

OpenStudy (frostbite):

In addition to that we notice that: \[\frac{ 1 }{ 2 }n(Mg)=n(O _{2})\]

OpenStudy (anonymous):

That formula is used to find the moles?

OpenStudy (frostbite):

Yea. you were wondering how to find the amount of mole... If I had to write the full equation it is going to look like this: \[V=\frac{ n(O _{2}) RT }{ p }\] Where n(O2) is given by: \[n(O _{2})=\frac{ 1 }{ 2 } \frac{ m(Mg) }{ M(Mg) }\] Insert into the equation: \[V=\frac{ \frac{ 1 }{ 2 } \frac{ m(Mg) }{ M(Mg) } RT }{ p }\]

OpenStudy (frostbite):

You see the idea?

OpenStudy (anonymous):

i get the idea, but where are you getting the 1/2 m(mg)/(M(mg))?

OpenStudy (frostbite):

Are you thinking about the 1/2?

OpenStudy (frostbite):

or the m(Mg)/M(mg)

OpenStudy (anonymous):

both

OpenStudy (frostbite):

Alright. the 1/2 come from the stoichiometry of the reaction as we say "2 moles of Mg react with 1 mole of dioxygen" so we have a 2:1 relationship. The m(Mg)/M(Mg) Come from the equation that say: \[n=\frac{ m }{ M }\] n = the amount of moles. m = the mass. M = the molecular mass.

OpenStudy (anonymous):

How do you figure out the ratio, is that the 2 in the equation?

OpenStudy (anonymous):

what is the significance of the 2 in front of Mg in the equation?

OpenStudy (anonymous):

ohh okay, so you assume a 1 in front of the O2 and that gives you the ratio.

OpenStudy (frostbite):

Exactly. If there is no number in front it is just 1... same we know from math 1 * x = x

OpenStudy (anonymous):

correct. but we know the molar mass, so what else is needed to figure out what the moles are?

OpenStudy (anonymous):

i came up with 25g X 1 mol ---- ----- 1 ? I don't get which molar mass i would decide to use, the oxygen or the magnesium?

OpenStudy (anonymous):

@Frostbite

OpenStudy (frostbite):

You are going to be using the molecular mass for magnesium in order to determine the amount of moles of oxygen.

OpenStudy (frostbite):

Btw I think you close to get of tack. Just put into the equations I set for you.

OpenStudy (frostbite):

out of track*

OpenStudy (anonymous):

ok so. to figure out the moles, the equation i have set is 1 6.022 X 10^23 25 X ----- X ------------- 24,305 1

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

okay, i got 1.03 moles of magnesium. Do i need the moles of oxygen as well? and do i add them together?

OpenStudy (frostbite):

No you don't need to add them together... what we are interested in the amount of moles of oxygen. So now when you have the amount of moles for magnesium you divide it with 2 and you then have the amount of moles of oxygen.

OpenStudy (anonymous):

I got the answer, thank you so much for your help!!(: